### v1

# Checklist Engineering Empowers Multilingual LLM Judges

Abstract
Automated text evaluation has long been a central issue in Natural Language Processing (NLP). Recently, the field has shifted toward using Large Language Models (LLMs) as evaluators—a trend known as the LLM-as-a-Judge paradigm. While promising and easily adaptable across tasks, this approach has seen limited exploration in multilingual contexts. Existing multilingual studies often rely on proprietary models or require extensive training data for fine-tuning, raising concerns about cost, time, and efficiency. In this paper, we propose Checklist Engineering based LLM-as-a-Judge (CE-Judge), a training-free framework that uses checklist intuition for multilingual evaluation with an open-source model. Experiments across multiple languages and three benchmark datasets, under both pointwise and pairwise settings, show that our method generally surpasses the baselines and performs on par with the GPT-4o model.111The code implementation is accessible at https://github.com/mghiasvand1/CE-Judge.
1 Introduction
Evaluation is one of the most important tasks in Natural Language Processing (NLP) for assessing a model’s performance on a specific task. Automating this process is highly beneficial and has been considered since the early stages. Traditional evaluation metrics (Papineni et al., 2002) have some drawbacks, such as the necessity of reference answers and a lack of interpretability, which has led to a paradigm shift toward developing Large Language Model (LLM) evaluators, referred to as LLM-as-a-Judge (Gu et al., 2024; Li et al., 2024). These models are also capable of evaluating long-form LLM generations in either a pointwise or pairwise format—meaning grading a single response or selecting the better response out of two, respectively. Some of the advantages of this approach include high adaptability (Bavaresco et al., ), low inference time, and the non-necessity for the evaluated LLM to be active during evaluation (i.e., it does not need to generate further responses). Despite significant efforts to make LLMs multilingual, extending LLM-as-a-Judge to multilingual configurations has received relatively little attention. Although current multilingual LLM judges (Pombal et al., 2025; Doddapaneni et al., 2024b) perform well, their main limitation is their reliance on proprietary models or the need for a large amount of real or synthetic data to fine-tune a capable evaluator, raising concerns regarding cost, time, and efficiency.
Meanwhile, checklists as interpretable evaluation tools (Doddapaneni et al., 2024a; Cook et al., ) are gaining traction for their transparency and structure, although their application to multilingual evaluation remains relatively underexplored. Most existing approaches also lack robust support for pairwise evaluation. For instance, (Wei et al., ) suggests to handle the pairwise setting by selecting the higher-graded response based on independent pointwise scores, which fails to capture the nuanced comparative superiority between responses. In this work, we present CE-Judge, an LLM-as-a-Judge framework that builds and uses engineered checklists for evaluation. It supports multilingual evaluation and both pointwise and pairwise modes. Our pipeline follows a three-stage process: the first two stages generate broad and dynamic checklist items, and the third applies them for judgment. Notably, even when using a lightweight, open-source model without any fine-tuning, our method demonstrates strong performance across different evaluation scenarios.
2 Related Works
2.1 LLM-as-a-Judge
We begin with Zheng et al. (2023), which uses the generative capabilities of LLMs to act as evaluators. These can be grouped into two types: prompt-based and fine-tuned evaluators. For the former, Li et al. (2025) adopts a ”decompose and aggregate” strategy, identifying weighted evaluation aspects and combining their scores to assess candidate responses. For the latter, which has recently gained momentum, Kim et al. (2024b) is a representative work that trains evaluators using large-scale synthetic data in both pointwise and pairwise setups, incorporating a weight-merging technique. In multilingual settings, consistency remains limited, as noted by Fu & Liu (2025). Among the few multilingual methods, Doddapaneni et al. (2024b) translates between languages to anchor outputs in English for consistent scoring, while Pombal et al. (2025) follows Kim et al. (2024b) to generate multilingual evaluation data and fine-tune models accordingly.
2.2 Checklist-based Evaluators
Several works have explored checklist-based evaluation. RocketEval (Wei et al., ) generates binary checklist items, then optimizes or reweights them to produce final scores. TICK (Cook et al., ) uses instructions to generate checklists, which LLMs use for self-improvement. CheckEval (Lee et al., 2024) defines high-level criteria, then decomposes, diversifies, and filters them to form evaluation checklists. FBI (Doddapaneni et al., 2024a) employs checklists for meta-evaluation to assess evaluator LLMs. Unlike these works, our framework introduces a novel architecture and, notably: (1) extends to multilingual settings; (2) supports pairwise evaluation beyond pointwise framing; and (3) uniquely incorporates broadness, descriptiveness, dynamism, and answer-mentioning in a unified manner.
3 CE-Judge pipeline
| Model | MMEval (Reasoning) | Avg. | ||||||||||
| en | de | fr | es | ru | zh | bn | ja | th | te | sw | ||
| Proprietary Models | ||||||||||||
| GPT-4o | 0.79 | 0.79 | 0.78 | 0.79 | 0.76 | 0.78 | 0.84 | 0.80 | 0.79 | 0.87 | 0.80 | 0.79 |
[…]
| Qwen2.5-7B-Instruct | ||||||||||||
| Hercule 7B | ||||||||||||
| M-Prometheus 7B | ||||||||||||
[…]
| Prometheus 2 8x7B | ||||||||||||
| M-Prometheus 14B | ||||||||||||
[…]
| CE-Judge | 0.77 | 0.81 | 0.77 | 0.72 | 0.78 | 0.77 | 0.75 | 0.84 | 0.78 | 0.76 | 0.78 | 0.77 |
We present our training-free, efficient evaluation framework (Figure 1), consisting of three steps. The pipeline builds an engineered checklist through level-by-level multilingual understanding, integrates input-output linkage for dynamism, and enhances LLM evaluator decisions. All LLM generations are forced in English to leverage its strong performance in that language.
3.1 Concepts generation
Considering an instruction as input—replaced by the source text in the translation evaluation task—and a corresponding response, we pass each separately to the LLM along with the prompt in A.1 for concept generation. This generation aims to produce an abstract-level text that represents the skeleton of the corresponding text.
3.2 Checklist generation
Next, we translate both the instruction and response texts into English. Using the translated response, the concepts generated from the instruction (from the previous step), and the prompt in A.2, we generate a checklist following the “response to instruction” direction. Likewise, we use the translated instruction and the response’s concepts to generate a checklist for the “instruction to response” direction. This dual approach aims to blind each side once, broadening checklist coverage and enhancing awareness of both sides’ content, rather than relying on a standard checklist with limited, predefined criteria. In this step, we also avoid prejudgment and ask the model to generate more descriptive items, going beyond simple binary questions.
3.3 Judgment
The final step is judgment. First, the two checklists from the previous steps are concatenated into a unified checklist. It’s important to note that the entire process described so far is for pointwise evaluation. For pairwise evaluation, the process remains the same, except that the previous two steps are applied to two candidate responses instead of one. As a result, we obtain two checklists, one for each candidate. In this step, we provide the untranslated versions of the instruction and response(s) (to avoid translation bias), along with the checklist(s) and the prompt template in A.3. The LLM is then asked to answer a subset of key checklist items and generate evaluation feedback. Unlike prior works where checklist items are marked with ticks, crosses, or weighted scores, here the model exercises discretion in its judgments, and the final evaluation is left to the model’s decision.
4 Experiments
| Model | MMEval (Chat) | Avg. | ||||||
| en | de | fr | es | ca | ru | zh | ||
| Proprietary Models | ||||||||
| GPT-4o | 0.72 | 0.73 | 0.78 | 0.73 | ||||
[…]
| Qwen2.5-7B-Instruct | 0.69 | 0.75 | 0.78 | 0.85 | ||||
| Hercule 7B | ||||||||
| M-Prometheus 7B | ||||||||
[…]
| Prometheus 2 8x7B | 0.72 | 0.77 | ||||||
| M-Prometheus 14B | 0.72 | |||||||
[…]
| CE-Judge | 0.69 | 0.73 | 0.77 | 0.82 | 0.77 | 0.87 | 0.75 | |
4.1 Experimental setup
In this work, we used the open-source Qwen2.5-7B-Instruct model (Yang et al., 2024) as the backbone LLM, accessed freely via the Novita API222https://novita.ai/. The hyperparameters “temperature”, “top_p”, and “seed” were set to 0, 1, and 42, respectively, to ensure reproducibility. For translation, we employed the free Google Translate API available through the deep-translator Python package333https://deep-translator.readthedocs.io/en/latest/README.html.
4.2 Datasets
Since our method is training-free, all datasets are used solely for testing. We evaluated our framework in both pointwise and pairwise settings. For pointwise evaluation, we used the student-annotated subset of the LitEval (Zhang et al., 2024), which contains source–target literary translations for four language pairs with human ratings from 1 to 7. For pairwise evaluation, we employed the reasoning and chat subsets of the MM-Eval dataset (Son et al., 2024), covering 11 and 7 languages, respectively. Each input consists of a reasoning question or chat history, with the task being to choose the better of two candidate responses.
4.3 Evaluation Metrics
| Model | LitEval | Avg. | |||
| deen | ende | enzh | dezh | ||
| Proprietary Models | |||||
| GPT-4o | |||||
[…]
| Qwen2.5-7B-Instruct | |||||
| Hercule 7B | |||||
| M-Prometheus 7B | 0.53 | 0.54 | 0.43 | ||
[…]
| Prometheus 2 8x7B | |||||
| M-Prometheus 14B | 0.29 | 0.57 | 0.48 | 0.56 | 0.47 |
[…]
| CE-Judge | 0.28 | 0.49 | |||
To evaluate our CE-Judge framework in pointwise mode, we used accuracy—defined as the number of correct predictions over the total number of samples. For the pairwise setting, we measured performance using Kendall’s Tau correlation coefficient (Kendall, 1938), which assesses the agreement between our model’s rankings and human judgments.
4.4 Baselines
We compare our framework with three types of models. The first includes proprietary models like GPT-4o. The second is Qwen2.5-7B-Instruct, a strong multilingual open-source LLM that is instruction-tuned from a pretrained model without further fine-tuning. The third category consists of models explicitly trained as evaluators, such as Prometheus 2 (Kim et al., 2024b), Hercule (Doddapaneni et al., 2024b), and M-Prometheus (Pombal et al., 2025), as discussed in Subsection 2.1.
5 Results
We evaluate CE-Judge on multilingual reasoning, chat, and literary translation tasks against proprietary and open-source baselines, including the fine-tuned M-Prometheus.
In the reasoning task (Table 3), CE-Judge achieves the highest average (0.77) among open-source models, outperforming several large fine-tuned evaluators like M-Prometheus 14B (0.70). Despite using only 7B parameters and no training, it performs close to GPT-4o (0.79), excelling in both high- and low-resource languages.
In chat evaluation (Table 4), CE-Judge leads again with 0.75 average accuracy, surpassing GPT-4o (0.73) and significantly outperforming M-Prometheus 7B (0.62).
For literary translation (Table 4.3), which requires fine-grained linguistic and stylistic insight, CE-Judge attains an average Kendall correlation (0.38) comparable to GPT-4o and much better than its backbone model (0.17). It slightly trails M-Prometheus 7B (0.43), which benefited from fine-tuning on large supervised multilingual data.
6 Conclusion
In this work, we introduce CE-Judge, a novel and straightforward checklist-based framework for multilingual LLM-as-a-Judge that is training-free and built on an open-source model. By leveraging dynamic, broad, and flexible checklist items, CE-Judge supports both pointwise and pairwise evaluations across diverse languages. Experiments on multiple multilingual benchmarks show that CE-Judge not only generally outperforms open-source baselines but also performs comparably to GPT-4o. These results highlight the promise of structured, dynamic evaluation techniques for improving the reliability and interpretability of LLM judgment, particularly in multilingual contexts, for more consistent performance.
7 Future directions
Despite its strong results and training-free design, our framework has several limitations to address in future work. First, crafting effective prompts for each of the three steps per task can be time-consuming, so developing an automatic, adaptable prompt generation module would be beneficial. Second, our method relies solely on LLM generation, which may suffer from misalignment between training objectives and robust text generation. Incorporating internal LLM representations, as shown by Sheng et al. (2024), could capture more accurate implicit knowledge. Finally, our framework’s flexibility suggests potential extensions as a plug-and-play method or adaptations to other evaluation strategies, such as interview-based evaluation (Kim et al., 2024a).
Appendix A Prompt templates
A.1 Concepts Generation Prompts
The prompts for this step, across all three datasets, are shown in 2, and the “[INPUT]” placeholder must be replaced with the text from which we want to extract concepts, such as an instruction, response, etc.
A.2 Checklist Generation Prompts
Figures 3, 4, and 5 show checklist generation prompts for Liteval, MM-Eval (Reasoning), and MM-Eval (Chat), respectively. Each figure consists of two prompts indicating the checklist creation direction. Note that the “[CONCEPTS]” placeholder must be replaced with the concepts generated in the previous step.
A.3 Judgment Prompts
We only use system prompts from this section, which are shown in 6: one for the Liteval dataset and another for the MM-Eval datasets. Figure 7 presents the prompt template for the Liteval dataset, while 8 shows the prompts for the two MM-Eval datasets. In these prompts, the placeholders clearly indicate what should replace them. Importantly, to demonstrate the flexibility of our framework, we also use a scoring guide for the pointwise assessment to help our judge LLM perform a more accurate evaluation.